OpenStudy (anonymous):

Use the distributive property to simplify the expression below: The quantity of 5 times x minus 25 all divided by 5

OpenStudy (cwrw238):

5( x - 25) -------- 5

OpenStudy (cwrw238):

that it?

OpenStudy (anonymous):

yes is the answer x-5

hero (hero):

Hmmm

hero (hero):

The fives cancel, no?

OpenStudy (cwrw238):

no its x-25 the 5's cancel out

OpenStudy (anonymous):

see by using bodmas we have (5x-25)/5 and hence its x-5

hero (hero):

If that's what it is, then @owenmisa123 misinterpreted @cwrw238's inquiry

OpenStudy (anonymous):

it happens random human errors...........

hero (hero):

Assuming @harsh314 is right: \[\frac{5x-25}{5} = \frac{5(x-5)}{5} = \frac{5}{5}(x-5) = 1(x-5) = x-5\]

OpenStudy (anonymous):

no not like that use distributive property \[\frac{ 5x-25 }{ 5 }=\frac{ 5x }{ 5 }-\frac{ 25 }{ 5 }=x-5\]

hero (hero):

@harsh314 What I posted IS distributive property. What you posted is called splitting the fraction.

OpenStudy (anonymous):

oh i have distributed the fraction 1/5 in the given polynomial but u have done a corollary of distributive by taking out 5 common actually u have not praciically distributed any number..............

hero (hero):

Distributive Property: ab - ac = a(b-c) 5x - 25 = 5(x-5)

hero (hero):

That's the distributive property. It's a PROPERTY. You don't have to distribute anything to use it

OpenStudy (anonymous):

no thats a corollary its a(b-c)= ab-ac

hero (hero):

The equal sign suggests that you can write the distributive property in reverse order and it is still the same.

OpenStudy (anonymous):

yes but plz dont mind but still the distributive word has a meaning

hero (hero):

Yes, but you're ignoring the fact that it says use "distributive PROPERTY"

OpenStudy (anonymous):

and it does not say use a corollary of distributive property neither have you distributed any of the quantity that has been given